OpenStudy (anonymous):

Can someone explain how to solve \[3^{-1/3}\times27^{-1/2}\] Thanks! :)

OpenStudy (mohsen):

\[3^{-1/3}\times3^{3(-1/2)}\] \[3^{-1/3+(-3/2)}\] \[3^{-1/3-3/2}\] \[3^{(-2-9)/6}\] \[3^{-11/6}\]

OpenStudy (anonymous):

how do you simplify that further? :D

OpenStudy (mohsen):

\[27=3^{3}\] then solve the fraction...

OpenStudy (mohsen):

\[a ^{m} \times a ^{n}=a ^{m+n}\] the base must equal...

OpenStudy (anonymous):

i mean, how do you simplify \[3^{-11/6}\] further? because actually, i got that answer too. but the answer isn't in the choices given. thanks!

OpenStudy (mohsen):

what the choices?

OpenStudy (anonymous):

\[\sqrt[6]{3}\] \[\sqrt[6]{3}/9\] \[3\sqrt[6]{3}/2\] and \[3\sqrt[6]{3}\]

OpenStudy (anonymous):

\[3^{-\frac{11}{6}} =3^{-2}.3^{1/6}={\sqrt[6]{3} \over 9}\]

OpenStudy (mohsen):

the answer is \[\sqrt[6]{3}/9\] \[3^{-11/6}=\] \[\sqrt[6]{3^{-11}}\] \[\sqrt[6]{3^{1}/3^{12}}\] \[\sqrt[6]{3^{1}}/\sqrt[6]{12}\] \[\sqrt[6]{3}/3^{12(1/6)}\] \[\sqrt[6]{3}/3^{2}\] \[\sqrt[6]{3}/9\]

OpenStudy (anonymous):

Anwara, how is 3−116=3−2.31/6 ?

OpenStudy (mohsen):

wow,anwar.how do u do that?

OpenStudy (anonymous):

i meant, how did you get to the 2nd equation? and thanks, mohsen!

OpenStudy (anonymous):

Because \(-2+1/6= \frac{-12+1}{6}=-11/6\), which is the same exponent :)

OpenStudy (anonymous):

@Mohsen: do what?

OpenStudy (mohsen):

same question with phoebeca :) bt u answered it :) thanks

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

thanks guys!

OpenStudy (mohsen):

i'm a teacher but still impress your answer,anwar :)

OpenStudy (anonymous):

i agree!

OpenStudy (anonymous):

Thanks mohsen :)